OpenStudy (anonymous):

solve the proportion 5/3c = 1/6

pooja195 (pooja195):

\[\huge~\rm~\frac{ 5 }{ 3c }=\frac{ 1 }{ 6 }\] Use "cross multiplication"

OpenStudy (anonymous):

\[\frac{ 5 }{ 3}c*\frac{ 3 }{ 5 } = \frac{ 1 }{ 6 }*\frac{ 3 }{ 5 }\]

pooja195 (pooja195):

\[\huge~\rm~5 \times 6 = ?\] \[\huge~\rm~3c \times 1= ?\]

OpenStudy (anonymous):

Thanks! I was blanking on what to do now I remember!

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)